The parenthesis always confuses me can someone help me with this? Im not very good with fractions it says multiply
9/5 times(-7/5)
any help is appreciated thanks

Answer by rfer(16322)

9/5*-7/5=-63/25=-2 13/25
() are used just to help explain the problem.
()() just means to multiplt